OpenStudy (anonymous):

x^2-4x=1 solve by completing the square

OpenStudy (anonymous):

-4x/2= -2 -2^2=4 (x-2)^2=1+4 x-2=square 5 x= 2 + - square 5

myininaya (myininaya):

x^2-4x+(4/2)^2=1+(4/2)^2 (x-2)^2=1+4 x-2=+ or - sqrt(5) x=sqrt(5)+2,-sqrt(5)+2

OpenStudy (anonymous):

\[(x-2)^{2} = 1+4\] \[(x-2)=\sqrt{5}\] \[x= 2\pm \sqrt{5}\]